# Multi-Source Position and Direction-of-Arrival Estimation Based on Euclidean Distance Matrices

### v1

@@ III-B Multi-Source Position Estimation
Similarly as for single-source position estimation, we consider $C$ TDOA estimates for the microphones $m=2,\dots,M$ with $C\geq S$. For each combination of candidate TDOA estimates, the optimal value $\hat{\alpha}(q)\,,\;q=1,\dots,Q,$ is determined using (28). Instead of only considering the overall smallest value of the cost function at these solutions, as in the single-source case, we now consider the $S$ smallest values $J(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1}),\dots,J(\hat{\alpha}(\hat{q}_{2}),\hat{q}_{S})$ and their corresponding Gram matrices $\hat{\mathbf{G}}_{1},\dots,\hat{\mathbf{G}}_{S}$. For each source, the relative microphones and source positions matrix $\hat{\mathbf{P}}_{\textrm{r}s}$ is estimated based on the eigenvalue decomposition of the corresponding Gram matrix $\hat{\mathbf{G}}_{s}$, similarly to (30). Then, similarly to (33), the position of the $s$-th source is then estimated as
@@
with $\hat{\mathbf{R}}_{s}$ the orthogonal mapping matrix for the $s$-th source, computed using the left and right singular vectors $\hat{\mathbf{M}}_{\textrm{r}s}\mathbf{M}^{\textrm{T}}$.
Since for reverberant and noisy scenarios with multiple sources the peaks of the GCC-PHAT function may sometimes contain spurious peaks that don’t correspond to the true TDOAs, it is often beneficial to set $C>S$.
For an exemplary scenario with $M=6$ spatially distributed microphones, $S=2$ sources (at distances $d_{11}=0.95$ m and $d_{12}=2.46$ m from the reference microphone) and $C=2$ candidate TDOA estimates, Fig. 2(a) illustrates the cost function $J(\alpha,q)$ in (26) for all $Q=32$ combinations of candidate TDOA estimates, while Fig. 2(b) shows the sorted minimum cost function values $J(\hat{\alpha}(q),q)$. For this scenario, it can be observed that only 2 out of 32 cost functions exhibit clear minima. Moreover, it can be observed that the estimated distance variables $\hat{\alpha}_{1}$ and $\hat{\alpha}_{2}$ for these combinations of TDOA estimates correspond very closely to the distances $d_{11}$ and $d_{12}$.
Figure 2: (a) Cost functions $J(\alpha,q)$ for all combinations of TDOA estimates $q$, (b) Corresponding minimum cost function values.
@@ IV-A Single-Source DOA Estimation
We consider a compact microphone array and a source in the far field of the microphone array, assuming that the acoustic waves arriving at the microphones from the source can be approximated as planar waves (see Fig. 3 for an exemplary two-dimensional configuration). This assumption is valid when the distances between the source and the microphones are much larger than the inter-microphone distances. In Section IV-A1 we define a rank-reduced Gram matrix, assuming the TDOAs to be known. In Section IV-A2 we explain how to incorporate multiple TDOA estimates.
@@ IV-A1 Rank-Reduced Gram Matrix
Figure 3: Exemplary two-dimensional configuration, consisting of a compact microphone array with $M=4$ microphones and a far-field source at azimuth angle $\theta_{1}$. In the relative coordinate system, defined by the DOA vector $\mathbf{v}_{1}$, the relative distances between the $m$-th microphone and the centroid of the microphone array, in the direction of the source, are denoted by $\delta_{m}\,,\,\;\;m=1,\dots,M$.
@@
with $\theta_{1}$ and $\phi_{1}$ the azimuth and elevation angle, respectively. These angles can be computed from the DOA vector as
$\theta_{1}\;$ $=\;\textrm{tan}^{-1}\left(\frac{\mathbf{e}_{2}^{\textrm{T}}\mathbf{v}_{1}}{\mathbf{e}_{1}^{\textrm{T}}\mathbf{v}_{1}}\right)\;,$ (37)
$\phi_{1}\;$ $=\;\sin^{-1}\left(\mathbf{e}_{3}^{\textrm{T}}\mathbf{v}_{1}\right)\;.$ (38)
@@
Using this mapping matrix, the microphone positions in the relative coordinate system (39) can be written in the absolute coordinate system, i.e., the relative microphone positions matrix $\mathbf{M}_{\textrm{r}}=\mathbf{R}_{1}^{\textrm{T}}\mathbf{M}$. The $M$-dimensional relative coordinate vectors $\mathbf{x}_{\textrm{r}}$, $\mathbf{y}_{\textrm{r}}$, and $\mathbf{z}_{\textrm{r}}$ are defined as
@@
It should be realized that the coordinate vector $\mathbf{x}_{\textrm{r}}$ contains the relative distances between the microphones and the centroid of the microphone array in the direction of the source (see Fig. 3). Since the relative distance $\delta_{m}$ between the $m$-th microphone and the centroid of the microphone array is directly related to the TDOA $\tilde{\tau}_{m}=\mathbf{m}_{m}^{\textrm{T}}\mathbf{v}_{1}/\nu$ between the $m$-th microphone and the centroid of the microphone array as $\delta_{m}=-\nu\tilde{\tau}_{m}$, the coordinate vector $\mathbf{x}_{\textrm{r}}$ can be written as
@@
where $\tilde{\boldsymbol{\tau}}=\left[\tilde{\tau}_{1},\dots,\tilde{\tau}_{M}\right]^{\textrm{T}}$ denotes the vector of (centered) TDOAs, for which $\tilde{\boldsymbol{\tau}}^{\textrm{T}}\mathbf{1}_{M}=0$.
@@
Using (42) and (43), it can be easily seen that $\mathbf{G}_{MM}^{\;-}=\mathbf{y}_{\textrm{r}}\mathbf{y}_{\textrm{r}}^{\textrm{T}}+\mathbf{z}_{\textrm{r}}\mathbf{z}_{\textrm{r}}^{\textrm{T}}$, whose rank is at most 2 (i.e., P1P-1). Considering the eigenvalue decomposition
$\mathbf{G}_{MM}^{\;-}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\textrm{T}}\;,$ (45)
with the matrix $\mathbf{W}$ containing the eigenvectors and the diagonal matrix $\boldsymbol{\Sigma}$ containing the eigenvalues, this means that at most $2$ eigenvalues are positive, $\sigma_{1}\geq\sigma_{2}\geq 0$, while the other eigenvalues are equal to zero, $\sigma_{3}=\dots=\sigma_{M}=0$. By defining
@@
the rank-reduced Gram matrix can be written as $\mathbf{G}_{MM}^{\;-}=(\mathbf{M}_{\textrm{r}}^{\;-})^{\textrm{T}}\mathbf{M}_{\textrm{r}}^{\;-}$. Based on (14), the matrix $\mathbf{M}_{\textrm{r}}^{\;-}$ can be computed from (45) up to an arbitrary orthogonal transformation $\mathbf{R}_{\textrm{ar}}^{\;-}$, i.e.,
$\mathbf{M}_{\textrm{ar}}^{-}\;=\;\left[\operatorname{\rm{diag}}\left(\sqrt{\sigma_{1}},\sqrt{\sigma_{2}}\right)\,,\;\mathbf{0}_{2\times(M-2)}\right]\mathbf{W}^{\textrm{T}}\;,$ (47)
with $\mathbf{M}_{\textrm{r}}^{-}=\mathbf{R}_{\textrm{ar}}^{-}\mathbf{M}_{\textrm{ar}}^{-}$. Using (42) and (46), the relative microphone positions can be written as
@@
with $\mathbf{R}^{\prime}$ an orthogonal mapping matrix. Realizing that the canonical basis vector $\mathbf{e}_{1}$ of the absolute coordinate system is equal to the first column of the matrix $\mathbf{R}_{\textrm{ar}}$ defined in (48), i.e., $\mathbf{e}_{1}=\mathbf{R}_{\textrm{ar}}\mathbf{e}_{1}$, the DOA vector $\mathbf{v}_{1}$ in (39) can be written as
@@
This means that the DOA vector is equal to the first column of the matrix $\mathbf{R}^{\prime}$ mapping the matrix $\mathbf{M}_{\textrm{ar}}$ to the absolute microphone positions matrix $\mathbf{M}$, which can be determined by solving an orthogonal Procrustes problem (without explicitly needing to compute the mapping matrices $\mathbf{R}_{1}$ and $\mathbf{R}_{\textrm{ar}}^{-}$).
@@ IV-A2 TDOA Estimation and Selection
Computing the matrices $\mathbf{G}_{MM}^{\;-}$ in (44) and $\mathbf{M}_{\textrm{ar}}$ in (48) requires centered TDOAs $\tilde{\boldsymbol{\tau}}_{m}$, which are not available in practice. Similarly as in Section III-A2, we consider $C$ candidate TDOA estimates $\hat{\tilde{\boldsymbol{\tau}}}_{m}\;,$$m=2,\dots,M$, corresponding to the $C$ largest local maxima of the GCC-PHAT function, and consider the $Q=C^{M-1}$ combination vectors $\mathbf{c}(q)$ in (23). For each combination $q$ of candidate TDOA estimates, the centered TDOA between the $m$-th microphone and the centroid of the microphone array can be estimated as
@@ IV-B Multi-Source DOA Estimation
It should be noted that due to possible TDOA estimation errors the rank of $\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{1})$ is not guaranteed to be equal to P1P-1. For multiple sources, we consider the $S$ smallest cost function values $I(\hat{q}_{1}),\dots,I(\hat{q}_{S})$ and their corresponding Gram matrices $\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{s})$. Using the eigenvalue decomposition $\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{s})=\hat{\mathbf{W}}(\hat{q}_{s})\hat{\boldsymbol{\Sigma}}(\hat{q}_{s})\hat{\mathbf{W}}^{\textrm{T}}(\hat{q}_{s})$, the matrices $\mathbf{M}_{\textrm{ar}}^{-}$ in (47) and $\mathbf{M}_{\textrm{ar}}$ in (48) can be estimated for each source as
@@
The mapping matrix $\hat{\mathbf{R}}^{\prime}(\hat{q}_{s})$ between the matrix $\hat{\mathbf{M}}_{\textrm{ar}}(\hat{q}_{s})$ and the absolute microphone positions matrix $\mathbf{M}$ can be computed by solving an orthogonal Procrustes problem, similarly to (31) and (32). From these mapping matrices, the DOA vectors can then be estimated using (50) as $\hat{\mathbf{v}}_{s}=\hat{\mathbf{R}}^{\prime}(\hat{q}_{s})\;\mathbf{e}_{1}$.
For an exemplary scenario with $M=6$ closely spaced microphones, $S=2$ sources (at distances $d_{\textrm{c}1}=1$ m and $d_{\textrm{c}2}=2$ m from the centroid of the microphone array) and $C=2$ candidate TDOA estimates, Fig. 4 shows the sorted cost function values I(q)I({q}) for all $Q=32$ combinations of TDOA estimates. For this scenario, it can be observed that there is a clear difference between the two smallest values and the other values.
Figure 4: Cost function values $I(q)$ for all combinations of candidate TDOA estimates $q$.

### v2

@@ III-B Multi-Source Position Estimation
Similarly as for single-source position estimation, in the TDOA estimation stage, no explicit association between the candidate TDOA estimates and sources is performed. We consider $C$ TDOA estimates for the microphones $m=2,\dots,M$, with $C\geq S$. For each combination of candidate TDOA estimates, the optimal value $\hat{\alpha}(q)\,,\;q=1,\dots,Q,$ is determined using (24). Instead of only considering the overall smallest value of the cost function at these solutions, as in the single-source case, the candidate TDOA estimates are automatically associated to different sources by identifying the $S$ of $Q$ combinations with the smallest values $J(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1}),\dots,\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\hat{\alpha}(\hat{q}_{S}),\hat{q}_{S})\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ and their corresponding Gram matrices $\hat{\mathbf{G}}_{1},\dots,\hat{\mathbf{G}}_{S}$. For each source, the relative microphones and source positions matrix $\hat{\mathbf{P}}_{\textrm{r}s}$ is estimated based on the eigenvalue decomposition of the corresponding Gram matrix $\hat{\mathbf{G}}_{s}$, similarly to (26). Then, similarly to (29), the position of the $s$-th source is estimated as
@@
with $\hat{\mathbf{R}}_{s}$ the orthogonal mapping matrix for the $s$-th source, computed using the left and right singular vectors of $\hat{\mathbf{M}}_{\textrm{r}s}\mathbf{M}^{\textrm{T}}$.
Since for reverberant and noisy scenarios with multiple sources the peaks of the GCC-PHAT functions may contain spurious peaks that don’t correspond to the true TDOAs, it may be beneficial to set $C>S$, particularly at lower SNRs, higher reverberation times, or for larger microphone spacings (see experimental results in Sections VI-E and VI-F). This, however, comes at the cost of increased computational complexity (see Section VI-D). In multi-source scenarios, where the GCC-PHAT function has an even higher probability of containing spurious peaks, it is even more beneficial than for single-source scenarios to set $C>S$. In fact, the proposed EDM-based method can be considered as a peak-to-source assignment method aiming at identifying the combinations of candidate TDOA estimates which correspond to the most feasible 3D geometry of sources and microphones. The EDM-based position estimation method is summarized in Algorithm 1 (supplementary material).
For an exemplary scenario, Fig. 5(a) (supplementary material) illustrates the cost function $J(\alpha,q)$ in (22) for $M=6$ spatially distributed microphones and $S=2$ sources, while Fig. 5(b) (supplementary material) shows the sorted minimum cost function values $J(\hat{\alpha}(q),q)$. For this scenario, it can be observed that only 2 out of 32 cost functions exhibit clear minima. Moreover, it can be observed that the estimated distance variables $\hat{\alpha}_{1}$ and $\hat{\alpha}_{2}$ for these combinations of TDOA estimates correspond very closely to the distances $d_{11}$ and $d_{12}$.
@@ IV-A Single-Source DOA Estimation
We consider a compact microphone array and a source in the far field of the microphone array, assuming that the acoustic waves arriving at the microphones from the source can be approximated as planar waves. (see Fig. 2 for an exemplary two-dimensional configuration). This assumption is valid when the distances between the source and the microphones are much larger than the inter-microphone distances. In Section IV-A1 we define a rank-reduced Gram matrix, assuming the TDOAs to be known. In Section IV-A2 we explain how to incorporate multiple TDOA estimates.
@@ IV-A1 Rank-Reduced Gram Matrix
Figure 2: Exemplary two-dimensional configuration, consisting of a compact array with $M=4$ microphones and a far-field source at azimuth angle $\theta_{1}$. In the relative coordinate system, defined by the DOA vector $\mathbf{v}_{1}$, the relative distances between the $m$-th microphone and the centroid of the microphone array, in the direction of the source, are denoted by $\delta_{m}\,,\,\;\;m=1,\dots,M$.
@@
with $\theta_{1}$ and $\phi_{1}$ the azimuth and elevation angle, respectively. These angles can be computed from the DOA vector as $\theta_{1}\;=\;\textrm{tan}^{-1}\left(\frac{\mathbf{e}_{2}^{\textrm{T}}\mathbf{v}_{1}}{\mathbf{e}_{1}^{\textrm{T}}\mathbf{v}_{1}}\right)\;,$ $\phi_{1}\;=\;\sin^{-1}\left(\mathbf{e}_{3}^{\textrm{T}}\mathbf{v}_{1}\right)\;.$ (33)
@@
The transpose of this matrix maps the microphone positions to the relative microphone positions as in (2), i.e., $\mathbf{M}_{\textrm{r}}=\mathbf{R}_{1}^{\textrm{T}}\mathbf{M}$. Equivalently, the matrix $\mathbf{M}_{\textrm{r}}$ describes the same microphone positions with respect to the absolute coordinate system as the matrix $\mathbf{M}$ describes with respect to the relative coordinate system. The $M$-dimensional relative coordinate vectors $\mathbf{x}_{\textrm{r}}$, $\mathbf{y}_{\textrm{r}}$, and $\mathbf{z}_{\textrm{r}}$ are defined as
@@
It should be realized that the coordinate vector $\mathbf{x}_{\textrm{r}}$ contains the relative distances between the microphones and the centroid of the microphone array in the direction of the source. (see Fig. 2). Since the relative distance $\delta_{m}$ between the $m$-th microphone and the centroid of the microphone array is directly related to the TDOA $\tilde{\tau}_{m}=\mathbf{m}_{m}^{\textrm{T}}\mathbf{v}_{1}/\nu$ between the $m$-th microphone and the centroid of the microphone array as $\delta_{m}=-\nu\tilde{\tau}_{m}$, the coordinate vector $\mathbf{x}_{\textrm{r}}$ can be written as
@@
where $\tilde{\boldsymbol{\tau}}=\left[\tilde{\tau}_{1},\dots,\tilde{\tau}_{M}\right]^{\textrm{T}}$ denotes the vector of (centered) TDOAs, for which $\tilde{\boldsymbol{\tau}}^{\textrm{T}}\mathbf{1}_{M}=0$. As will be shown in the next subsection, these centered TDOAs can be estimated despite there not necessarily being a microphone at the centroid of the microphone array.
@@
Using (37) and (38), it can be easily seen that $\mathbf{G}_{MM}^{\;-}=\mathbf{y}_{\textrm{r}}\mathbf{y}_{\textrm{r}}^{\textrm{T}}+\mathbf{z}_{\textrm{r}}\mathbf{z}_{\textrm{r}}^{\textrm{T}}$, whose rank is at most 2 (i.e., P1P-1). Considering the eigenvalue decomposition $\mathbf{G}_{MM}^{\;-}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\textrm{T}}$ with the matrix $\mathbf{W}$ containing the eigenvectors and the diagonal matrix $\boldsymbol{\Sigma}$ containing the eigenvalues, this means that at most $2$ eigenvalues are positive, $\sigma_{1}\geq\sigma_{2}\geq 0$, while the other eigenvalues are equal to zero, $\sigma_{3}=\dots=\sigma_{M}=0$. By defining
@@
the rank-reduced Gram matrix can be written as
$\mathbf{G}_{MM}^{\;-}=(\mathbf{M}_{\textrm{r}}^{\;-})^{\textrm{T}}\mathbf{M}_{\textrm{r}}^{\;-}\;.$ (41)
As explained in Section II, a Gram matrix is only defined up to an arbitrary orthogonal transformation. This means that the rank-reduced Gram matrix $\mathbf{G}_{MM}^{\;-}$ can also be written as $\mathbf{G}_{MM}^{\;-}=\left(\mathbf{M}_{\textrm{ar}}^{-}\right)^{\textrm{T}}\left(\mathbf{M}_{\textrm{ar}}^{-}\right)$, with $\mathbf{M}_{\textrm{ar}}^{-}=\left(\mathbf{R}_{\textrm{ar}}^{-}\right)^{\textrm{T}}\mathbf{M}_{\textrm{r}}^{-}$, and $\mathbf{R}_{\textrm{ar}}^{-}$ an arbitrary orthogonal matrix. Based on (10), the matrix $\mathbf{M}_{\textrm{ar}}^{\;-}$ can be computed as
$\mathbf{M}_{\textrm{ar}}^{-}\;=\;\left[\operatorname{\rm{diag}}\left(\sqrt{\sigma_{1}},\sqrt{\sigma_{2}}\right)\,,\;\mathbf{0}_{2\times(M-2)}\right]\mathbf{W}^{\textrm{T}}\;\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ (42)
Using (37) and (40), the relative microphone positions can be written as
@@
with $\mathbf{R}^{\prime}$ an orthogonal matrix mapping the matrix $\mathbf{M}_{\textrm{ar}}$ to the absolute microphone positions matrix $\mathbf{M}$, which can be determined by solving an orthogonal Procrustes problem. Realizing that the canonical basis vector $\mathbf{e}_{1}$ of the absolute coordinate system is equal to the first column of the matrix $\mathbf{R}_{\textrm{ar}}$ defined in (43), i.e., $\mathbf{e}_{1}=\mathbf{R}_{\textrm{ar}}\mathbf{e}_{1}$, the DOA vector $\mathbf{v}_{1}$ in (34) can be written as
@@
This means that the DOA vector is equal to the first column of the mapping matrix $\mathbf{R}^{\prime}$ and does not depend on the arbitrary orthogonal matrix $\mathbf{R}^{-}_{\textrm{ar}}$.
@@ IV-A2 TDOA Estimation and Selection
Computing the matrices $\mathbf{G}_{MM}^{\;-}$ in (39) and $\mathbf{M}_{\textrm{ar}}$ in (43) requires centered TDOAs $\tilde{\boldsymbol{\tau}}_{m}$, which are not available in practice. Similarly as in Section III-A2, we consider $C$ candidate TDOA estimates $\hat{\tilde{\boldsymbol{\tau}}}_{m}\;,$$m=2,\dots,M$, corresponding to the $C$ largest local maxima of the GCC-PHAT function, and consider the $Q=C^{M-1}$ combination vectors $\mathbf{c}(q)$ in (19). Similarly as for position estimation, it should be noted that the number of combinations $Q$ increases exponentially with $M$. For each combination $q$ of candidate TDOA estimates, the centered TDOA between the $m$-th microphone and the centroid of the microphone array can be estimated as
@@ IV-B Multi-Source DOA Estimation
It should be noted that due to possible TDOA estimation errors the rank of $\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{1})$ is not guaranteed to be equal to P1P-1. For multiple sources, we consider the $S$ smallest cost function values $I(\hat{q}_{1}),\dots,I(\hat{q}_{S})$ and their corresponding Gram matrices $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{1}),\dots,\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{S})\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$. Using the eigenvalue decomposition $\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{s})=\hat{\mathbf{W}}(\hat{q}_{s})\hat{\boldsymbol{\Sigma}}(\hat{q}_{s})\hat{\mathbf{W}}^{\textrm{T}}(\hat{q}_{s})$, the matrices $\mathbf{M}_{\textrm{ar}}^{-}$ in (42) and $\mathbf{M}_{\textrm{ar}}$ in (43) can be estimated for each source as
@@
The mapping matrix $\hat{\mathbf{R}}^{\prime}(\hat{q}_{s})$ between the matrix $\hat{\mathbf{M}}_{\textrm{ar}}(\hat{q}_{s})$ and the absolute microphone positions matrix $\mathbf{M}$ can be computed by solving an orthogonal Procrustes problem, similarly to (27) and (28). From these mapping matrices, the DOA vectors can then be estimated using (45) as $\hat{\mathbf{v}}_{s}=\hat{\mathbf{R}}^{\prime}(\hat{q}_{s})\;\mathbf{e}_{1}$. No additional regularization beyond the eigenvalue truncation in (50) is applied for solving the orthogonal Procrustes problem. The EDM-based DOA estimation method is summarized in Algorithm 2 (supplementary material).
For an exemplary scenario, Fig. 6 (supplementary material) shows the sorted cost function values I(q)I({q}) for $M=6$ closely spaced microphones and $S=2$ sources. For this scenario, it can be observed that there is a clear difference between the two smallest values and the other values.